In 1993, Nash Company completed the construction of a building at a cost of $2,040,000 and first occupied it in January 1994. It was estimated that the building will have a useful life of 40 years and a salvage value of $59,200 at the end of that time. Early in 2004, an addition to the building was constructed at a cost of $510,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $20,400. In 2022, it is determined that the probable life of the building and addition will extend to the end of 2053, or 20 years beyond the original estimate.

Compute the annual depreciation to be charged, beginning with 2022. (Round answer to 0 decimal places, e.g. 45,892.)
Answer #1

In 1993, Nash Company completed the construction of building, see the below information

Cost of building in 1994 = $2,040,000

Estimated Usage life = 40 years

Salvage Value = $59,200

In 2004 :

Additional cost of building = $510,000

Additional usage life = 30 years

Estimated Salvage value = $20,400

Depreciation = (Cost of building - Salvage Value)/usage life

Depreciation = (2,040,000 - 59,200)/40

Depreciation = 49,520 per year

From 1994 to 2003 total 10 years so,

cost of building in 2004 =2,040,000-(49,520*10)

cost of building in 2004 = $1,544,800

In 2004, with additional cost

cost = 1,544,800+510,000 .

= 2,054,800

Salvage Value = $79,600

Estimated life = (40-10)+30

Estimated Life = 70 years

From 2004,the estimated life would be till 2073th Year

but in 2022, the life of building will be extended to 2053, 20 years beyond estimated life.

If the life upto 2053,from 2004 the estimated life would be 50 years.

Annual depreciation from 2022 = (2,054,800 - 79,600)/50

= $39,504